# A Review on Deep Learning Autoencoder in the Design of Next-Generation Communication Systems

### v1

@@ X-B Optical Communications
[149] adaptability and robustness of quantum neural networks (QNNs) in representing complex classical problems hybrid quantum-classical AE using QNN standard encoded radio signals, AWGN and Rayleigh fading Superior performance with reduced parameters and maintained accuracy, particularly with quantum encoders Framework optimization for near-term quantum devices, considering the computational challenges of complex modulations
@@
In the field of optical fibre communication, Uhlemann et al. [129] explored the optimisation of coherent optical communication systems, particularly under the influence of CD and Kerr nonlinearity (KNL), using deep learning models to mitigate dispersion and interference. The fibre channel model is implemented using the Split-Step Fourier Method (SSFM). Alternatively, the trade-off between amplification noise and nonlinear fibre distortions is addressed in [130]. The aim is to perform GCS optimisation that mitigates both amplification noise and nonlinear effects in fibre optic communication systems. Optimisation of transceiver designs has also been carried out in the short-reach optical fibre link, where intensity modulation with direct detection (IMDD) [131, 132]. The proposed AEs use artificial neural networks (ANNs) with the emphasis on the performance of feedforward and bidirectional recurrent neural networks (BRNNs) in mitigating dispersion and interference in the optical channel. A differentiable link has been modelled that accounts for CD and memory effects in the optical fibre channel.
Some studies focused on designing AEs with an accurate channel modelling and incorporating DSP techniques in coherent optical communication systems, demonstrating significant improvements over traditional methods. Some approaches explored the use of VAEs for blind equalisation and channel estimation. Lauinger et al. [133] demonstrated that VAEs can outperform traditional methods such as the constant modulus algorithm (CMA) in fixed and time-varying channels. A dispersive linear optical dual-polarisation channel is employed, along with an AWGN and intersymbol interference (ISI) time-varying channel. On the other hand, Song et al. [134] presented a novel frequency domain blind equalisation scheme using vector-quantised variational AEs (VQ-VAEs). It demonstrated that this approach can achieve performance similar to that of time-domain methods while reducing computational complexity, particularly in long-memory polarisation division multiplexing (PDM) channels. This implies that the VQ-VAE equaliser can overcome the slow convergence of traditional blind equalisation methods, such as CMA, especially for high-order modulation formats. Similarly, Lauinger et al. [135] focused on bootstrapping blind adaptive equalisers, such as CMA. Novel VAE-based equalisers demonstrated reliable convergence at these critical points for both standard and PCS formats. The demonstration used a fibre channel using a linear frequency domain channel matrix, assuming a time-invariant channel during the bootstrapping phase AWGN and first-order PMD. Those works used the evidence lower bound (ELBO) modified loss function to assess end-to-end learning of the proposed VAE approach. Other approaches allow for end-to-end training to enhance system performance in phase-noise effected channels. The works in [136, 137, 138] investigated the optimisation of both geometric and probabilistic constellation shaping within Wiener phase noise channels, consequently enhancing spectral efficiency and data throughput. In particular, the integration of carrier CPE into the optimisation process, which requires a differentiable approach of CPE and BPS to enable end-to-end optimisation.
In general, these studies collectively underscore the potential of deep learning in revolutionising optical communication systems by providing robust, efficient, and adaptive solutions to complex channel conditions.
X-C Semantic Communication
An AE-based semantic communication system (AESC) is a promising architecture to enhance the reliability of semantic communication in wireless communication. It emphasises the importance of semantic information over traditional data recovery methods, focussing on conveying meaning rather than just symbols. The potential of using conceptual spaces to model meaning geometrically is emphasised in [146], which addresses the limitations of existing black-box frameworks in semantic communication. A modified AE architecture is introduced that uses dense layers and convolution layers with the ReLU activation function. The CCE is considered to ensure that the classification vectors produced by the network are close to the true property vectors, along with a function, which is a reconstruction loss associated with the AE output, to penalise large distortion. On the other hand, Samarathunga et al. [147] presented an objective quality metric based on AE designed to quantify semantic noise in image transmission systems. They discussed the inadequacy of existing metrics in addressing semantic noise, which is crucial for the success of semantic communication systems. The proposed model demonstrated a high correlation (88%) with subjective quality assessments, indicating its effectiveness in evaluating semantic noise. A convolutional AE architecture is utilised to handle semantic meaning. The MSE between latent vector spaces of undistorted and distorted images is used to quantify semantic noise alone with BCE for end-to-end learning.
X-D Quantum Communication
Exploring the combination of quantum and classical approaches to enhance end-to-end communication systems is a promising area for further research. The authors in [149] and [150] presented a hybrid quantum-classical AE designed for end-to-end radio communication, leveraging Quantum Neural Networks (QNNs) to enhance performance in noisy environments. The works emphasised the adaptability and robustness of QNNs in representing complex classical problems, particularly improving robustness against signal degradation and meeting inference-time constraints through a generalised data re-uploading scheme. Hybrid Quantum-Classical AE (QC-AE) architecture systems are introduced, focussing on AWGN and Rayleigh fading channels. The QC-AE is trained symbol-wise using SCCE.
@@ X-D Quantum Communication
Rathi et al. [151] highlighted the utility of parametrised quantum circuits to optimise channel codes in classical, entanglement-assisted, and quantum communication scenarios. Strong performance is achieved with qubit-based frameworks in various quantum noise channels such as bit-flip, phase-flip, and depolarising models, with learnt capacities approaching theoretical limits. They used Hilbert space manipulation with loss functions tailored to classical and quantum communication capacities (cross-entropy for classical tasks, trace distance for quantum settings).
X-E AE Enhancements and Challenges
TABLE XIII: Comparison of AE Enhancements and Challenges
Ref. Main Contribution Model(s) Used Dataset/Data type/Channel model Best performance value Limitation
[21] BMD optimisation Dense layers (ReLU, sigmoid) SDRs, AWGN Significant gains Gaussian LLRs assumption
[28] Joint Tx/Rx optimisation SPSA AE, conv1D (Tanh) Random binary, AWGN Outperforms 64QAM Reliance on synthetic channels
[38] Optimal data representations Dense (ReLU), softmax Synthetic, AWGN, RTNs 4 dB in low-medium SNR Training at low SNR hinders learning
[40] Enhance communication systems Dense layers, embedding reg. Modulated signals, AWGN Reduced PAPR, close to conventional BER Trained at single $E_{b}/N_{0}$
[152] Improved data transmission Dense (ReLU), symbol-wise Simulated, stochastic AWGN BLER below $10{-3}$ at SNR ¿ 0 dB Evaluated only in AWGN
[153] High complexity reduction Convolutional, BNNs Random symbols, Rayleigh Storage: 1/7, Computation: 1/5 High complexity of current AE-based systems
[154] Phase shift compensation ANN, dense, bit-wise AWGN, real channel Effective compensation up to SNR 10 dB High computational complexity
[155] Receiver synchronisation Dense (ReLU), softmax Random messages, AWGN, CFO BLER 1 dB worse than GR DQPSK Challenge of training over actual channels
[156] Differentiable channel models Control layer, softmax One-hot encoder vectors, AWGN, RBF Faster convergence, stronger robustness Reliance on mathematical assumptions
[157] Model real-world channel effects Dense (ReLU), softmax Over-the-air signals, AWGN 99.8% classification accuracy Adversarial ML challenges, data adaptation
[158] end-to-end optimisation without channel knowledge VAE, symbol-wise, convolution Simulated, AWGN, Rayleigh, Rician Improved SER Further evaluation for 64PSK, 128PSK
[85] Design capacity-approaching codes Dense, symbol-wise Probabilistic channel model Improved decoding schemes Dependency on training, MI estimator instability
[149] Efficient data transmission Hybrid QC-AE, symbol-wise, Dense Standard encoded radio, AWGN, Rayleigh Superior performance, reduced parameters optimisation for near-term quantum devices
[150] optimise learning rate Hybrid AE, symbol-wise, QNN 2D encoded symbols, AWGN 40% accuracy improvement Inference time exceeds acceptable latency
[159] Simplify communication system Dense (ReLU), softmax Multipath, Additive random noise SER reduction to SNR 12 dB Did not converge at certain amplitude values
[160] optimise discrete distributions Tx: Dense (ReLU, linear), Rx: Dense (ReLU) AWGN, Rayleigh fading Close to capacity on AWGN Finding optimal marginal distribution is challenging
[161] optimise constellation for near Shannon capacity Dense (ReLU), softmax Shared communication, AWGN Improves SER over TDMA/FDMA computational complexity
[162] Relax interconnect bandwidth Dense, MSE MIMO Wireless Communication Interconnect bandwidth relaxed by up to 8 Computational complexity of decentralized processing
[163] optimisation in high-noise environments ResNet, BCE AWGN BER improvement from 11 dB to 6 dB Does not achieve better performance on large block lengths
[164] Effective coding schemes Dense (ReLU), Softmax Probabilistic channel model Effectiveness of proposed methodology Scaling challenges, MI estimation inaccuracy
[165] Investigate various channel conditions CNN, LSTM Random binary, AWGN, multipath BER can be reduced to acceptable range Initial performance of block scheme not sufficient
[166] Susceptibility of DNN AEs MPL AE, CNN AE AWGN, adversarial attack BLER significantly increases at PSR $-2$ dB Vulnerability to adversarial attacks
[168] High-reliability maritime communication DenseNet-MAE Random binary, AWGN, Rician fading BLER $10{-3}$ in Rician fading Training at low/high SNR leads to suboptimal generalisation
[169] Manage channel effects, interference Fully connected (ReLU, dropout) OFDM channel, AWGN Up to 36 dB interference suppression performance variation with antenna/interference configurations
[170] Generation of effective constellations Dense (ReLU) Simulated data, IEEE 802.11a Constant amplitude QPSK, resulting in a more rounded constellation of QPSQ QPSK improvement, 16/64QAM initial performance issues
X-E1 AE Enhancement
Some researches have investigated the enhancements of AE architectures. AEs models are based on employing one-hot vector representations, which yields, however, low data rates. Chen et al. [152] introduced a novel generalised data representation (GDR) scheme aimed at enhancing the data rate in DL-based communication systems. The demonstration revealed that the GDR can significantly improve channel capacity and BLERperformance. Various learning strategies for AEs, such as weight initialisation and hyperparameter selection, are explored in [39]. However, the proposed AE has been evaluated using the AWGN channel. It is crucial to address different channel impairments to identify the appropriate learning strategy that can effectively capture channel impairments. Furthermore, Tang et al. [171] focused on the lack of interpretability in AE models, especially in critical fields like communication. It demonstrated how a meta-model can approximate the AE using a generalised hypergeometric function, allowing for a clearer understanding of its operations and performance.
X-E2 High-Computational Complexity
several researches have covered the topic of high computational complexity and resource limitations of generally implementing ANNs and particularly AEs on FPGAs for communication systems. In [122] a low complexity architecture has been proposed to address the challenge of high complexity of the model in advanced channel-coded systems and higher-order modulation schemes. Alternatively, the study in [153] developed a Binary Neural Networks (BNN)-based AE that can operate effectively on hardware with limited computational resources. It used low-precision compression techniques, specifically one-bit quantisation. Moreover, the authors of [154] proposed a hybrid approach that combines the adaptability of AEs with the efficiency of conventional demapping algorithms. The model is evaluated on an abstract channel model for end-to-end training of the system like AWG), followed by retraining under real channel conditions by varying the phase-shift.
X-E3 Non-Differentiable and Model-Free Channel
The need for differentiable channel models for training neural networks in communication systems is crucial to enable gradient backpropagation. Due to real-world channels being non-differentiable, many researchers have investigated a way to guarantee global optimality during end-to-end training. Dorner et al. [155] proposed a two-step learning procedure based on transfer learning to effectively train the AE model. First, the AE is trained using a stochastic channel model that simulates the expected behaviour of the actual communication channel. Then, fine-tuning with real data is carried out after deployment on real hardware for over-the-air transmissions. This step is important because it helps the system adapt to real-world channel conditions, which may differ from the stochastic model used. Another solution is to add reinforcement learning (RL) to the training process. In [27], a novel learning algorithm is proposed that combines supervised training for the receiver with RL for the transmitter, allowing training without prior knowledge of the channel model. Furthermore, Xu et al. [78] used RL to estimate channel models. However, integrating RL with the AE architecture adds complexity to the training stage. One solution is to use the simultaneous perturbation stochastic approximation (SPSA) algorithm instead of backpropagation [28]. Finally, Hu et al. [156] proposed a Kalman-based AE framework that integrates Kalman filtering with DNNs. The framework utilises two types of Kalman filters: the extended Kalman filter (EKF) for differentiable channel models and the cubature kalman filter (CKF) for non-differentiable models. This dual approach allows robust performance across various channel conditions, including AWGN and rayleigh block splitting (RBF).
The main challenge of designing a communication system is accurately modelling real-world channel effects, which are often oversimplified in existing mathematical models. Therefore, the need for model-free approaches that can learn channel responses from data without relying on closed-form models is essential. Erpek et al. [157] introduced a GAN to approximate the channel response, allowing the modelling of arbitrary communication systems. This approach eliminates the need for predefined channel models, accommodating real-world hardware imperfections and nonlinearities complexities. The idea has been extended in [158], where RL and GAN are proposed to develop model-free communication frameworks. It emphasised the representation of information messages as packet hot vectors (PHV) inferred from latent random variables (LRVs), allowing for the transmission of only statistical parameters rather than exact bits, which is significantly improving performance compared to traditional methods. On the other hand, Letizia et al. [85] addressed the design of capacity-approaching codes for unknown channels, where the channel capacity is not predefined. The work presented a novel approach to AE training by integrating mutual information as a regularisation term in the cross-entropy. This method aims to enhance the learning of capacity-approaching codes that maximise mutual input-output information under power constraints, thereby addressing overfitting issues commonly associated with traditional training methods.
XI Future Research Areas
Future research can further advance the application of AEs in communication systems, leading to more efficient, reliable, and high-performance solutions. Therefore, we identify the following research directions.

### v2

@@ VI-D3 Background Knowledge Alignment
[135] Adaptability and robustness of QNNs in representing complex classical problems hybrid quantum-classical AE using QNN standard encoded radio signals, AWGN and Rayleigh fading Superior performance with reduced parameters and maintained accuracy, particularly with quantum encoders optimization for near-term quantum devices, considering complex modulations computational
VI-E Hybrid Quantum-Classical Communication
This emerging field combines quantum and classical techniques, which introduces unique challenges related to integrating these different paradigms, managing noise in quantum systems, and optimizing performance on near-term quantum hardware, as listed in Fig. 10.
Hybrid Quantum-Classical CommunicationSolutionsChallenges Quantum Dataset and Encoding: high-dimensional quantum datasets, adapting to different encoding schemes and scaling models System Integration and Optimization: high-performance near real-time, hyperparameter, adapting QNN architecture Hybrid quantum-classical AE architecture … active challenge area Hybrid quantum-classical AE architecture … active challenge area
Figure 10: Hybrid quantum-classical communication challenges and solutions
VI-E1 Quantum Dataset and Encoding
The integration of quantum machine learning techniques into communication systems introduces a new set of challenges in relation to quantum hardware limitations and performance variability. In experimental settings, factors such as fluctuating photon numbers, variable bin sizes, and measurement noise complicate the application of conventional dimensionality reduction tools like t-SNE. Notably, the sensitivity of t-SNE to hyperparameters often results in overlapping clusters, which reduces interpretability in high-dimensional quantum datasets [136]. Furthermore, the performance of quantum models degrades significantly when the quantum efficiency drops below 0.2, while saturation effects in classification accuracy are observed as bin sizes exceed optimal thresholds.
Although VAE models demonstrate robustness in such noisy quantum environments, they must contend with inherent losses and noise associated with real-world quantum experimental setups [136]. As such, there is a growing need to optimize these models for near-term quantum devices, which are constrained by limited qubit counts, decoherence, and gate infidelities [135]. Furthermore, adapting to different encoding schemes and scaling models effectively across diverse quantum architectures remain active areas of exploration. In response, recent work has introduced hybrid quantum-classical AE frameworks designed for end-to-end radio communication, effectively leveraging the representational capacity of QNNs to model channel behaviors and reconstruct signals with enhanced efficiency [134]. These systems demonstrate the viability of QNNs in channel autoencoder architectures, which offer potential gains in training speed, inference latency, and robustness against quantum noise [135]. Moreover, optimization of quantum circuit designs has been shown to reduce training overhead and improve runtime performance. This highlights a clear path forward for deploying QNNs in near-term quantum-enabled communication platforms.
To further support the development of QNNs for practical applications, more extensive studies are required to evaluate their performance under noisy channel conditions and to design architectures that meet stringent latency and resource constraints [135].
VI-E2 System Integration and Optimization
Despite recent progress in quantum machine learning (QML) for communication systems, several challenges remain unaddressed in the context of system-level integration and real-time performance. In particular, high-performance near real-time applications represent an area where further empirical investigation is required to assess practical viability under operational constraints [134]. Performance under complex channel noise conditions also requires a more rigorous study, especially as QNN models scale to broader deployment environments. Another persistent challenge lies in hyperparameter tuning during end-to-end system training, which plays a critical role in minimizing estimation errors and improving overall system generalizability [134]. Furthermore, adapting QNN architectures to diverse and noisy channel conditions presents both theoretical and implementation-level complexities, particularly when integrating data re-uploading mechanisms and tailoring core circuit structures to specific application demands [135]. The adaptability of these models to general encoding schemes (beyond standard modulation formats such as 4-QAM and 16-QAM) remains an important avenue for future research.
To address these issues, a hybrid quantum-classical AE architecture has been proposed for end-to-end radio communication, in which a quantum decoder is paired with a classical encoder [134]. This approach allows for the incorporation of quantum computation expressive power while mitigating the limitations of current quantum hardware. Notably, a generalized data re-uploading strategy has been introduced for qubit-based circuits to enable the system to satisfy stringent inference time constraints without compromising representational fidelity. Initial findings suggest that these hybrid architectures are promising candidates for communication over noisy channels.
@@ VI-E2 System Integration and Optimization
Rathi et al. [137] highlighted the utility of parametrized quantum circuits to optimize channel codes in classical, entanglement-assisted, and quantum communication scenarios. Strong performance is achieved with qubit-based frameworks in various quantum noise channels such as bit-flip, phase-flip, and depolarizing models, with learned capacities approaching theoretical limits. They used Hilbert space manipulation with loss functions tailored to classical and quantum communication capacities (cross-entropy for classical tasks, trace distance for quantum settings).
However, the performance of the hybrid AE in more complex channel noise scenarios requires additional study to validate its superiority over purely classical methods. Additionally, the is open challenge when adapting the model to higher numbers of qubits, as this requires adjustments to the QNN building blocks and necessitates a new choice of measurement observables, which can complicate the implementation.
VI-F Other Specific Challenges
Additional challenges associated with existing solutions are illustrated in Fig. 11. These pertain more broadly to the design of communication systems as a whole.
Other Specific Topics in Communication SystemsSolutionsChallenges Complexity and Computational Constraints: computation and storage demands in resource-constraint HW Training and Convergence: high-dimensional data, ensure convergence Data Representation and Rate: one-hot low spectral efficiency Data Scarcity: Collecting/Classifying RF signals, large labeled datasets Tensor decomposition, shallow network architectures, low-precision compression, binary NN, hybrid architecture … Segmented GAN architecture, Community-based AE, Two-step training Generalized data representation (GDR), … active challenge area VQ-VAE, deep denoising AE, semi-supervised classification, … open challenge
Figure 11: Other specific challenges and solutions
VI-F1 Complexity and Computational Constraints
Deploying data-driven NN transceivers in practical communication systems is challenged by significant computational and storage demands, especially on resource-constrained hardware [22]. The inherent complexity of AE-based architectures adversely impacts system scalability and real-time performance [21]. Training and inference in artificial neural networks (ANNs), particularly large-scale models, require substantial computational resources, limiting their applicability in low-power or embedded environments [22]. Additionally, the large parameter space of ANNs further restricts deployment due to storage limitations [22].
To address these challenges, several solutions have been proposed. Tensor decomposition techniques (such as CP, Tucker, TT, and TR) provide an effective means to reduce both computational complexity and memory requirements without compromising performance [22]. These techniques result in performance degradation (0.75dB to 1.3dB bit error rate degradation) that is directly related to the compression technique facilitated by tensor decomposition. In multi-user systems, shallow network architectures have been demonstrated to achieve satisfactory results while maintaining low complexity [21]. Furthermore, low-precision compression methods, such as one-bit quantization, enable efficient end-to-end learning by reducing computational burden [62]. Binary neural networks (BNNs) represent a promising direction, offering near-equivalent performance to full-precision models with significantly reduced complexity [62]. Although using low-precision compression methods and related architectures such as BNNs are presented as beneficial for complexity reduction, the challenge of storage and computation complexities can remain significant for end-to-end learning communication systems; this is less a limitation of the low-precision method as a technique for reduction. A hybrid architecture has also been proposed that combines the adaptability of AEs with the efficiency of conventional demapping, striking a balance between performance and computational feasibility [22]. The hybrid approaches can achieve optimized performance while maintaining hardware efficiency and addressing complexity [22].
VI-F2 Training and Convergence
Training NN for communication systems presents several challenges, particularly in the presence of high-dimensional data and channel uncertainty. As network size increases, especially with large input dimensions (denoted by $n$), the training process becomes susceptible to unreliable convergence and entrapment in local minima, which undermines the reliability of performance [29]. Additionally, training models based on inaccurate or incomplete channel representations can lead to suboptimal generalization and noticeable performance degradation when deployed on actual communication channels [62]. Ensuring convergence to an optimal solution across a diverse range of channel conditions remains a fundamental difficulty, exacerbated by the significantly increased complexity associated with training large-scale neural network architectures [94].
To address these challenges, several advanced solutions have been proposed. A segmented GAN architecture incorporating copula theory has been introduced to address challenges in training NNs, particularly when dealing with high-dimensional data and unknown channels [29]. A community-based AE architecture enables joint optimization of transmitter and receiver functions to mitigate the limitations of prior decoupled designs [38]. It is also used to tackle the challenge of ensuring that the learning procedure converges to an optimal solution for every channel. Practical end-to-end learning frameworks have also been developed, such as trainable point-to-point systems where both transmitter and receiver neural networks are optimized using real-world wireless channel data, which can improve both adaptability and robustness [14]. Furthermore, a two-step training paradigm based on transfer learning has been proposed to overcome difficulties associated with real-channel variability, which offers enhanced convergence and generalization across heterogeneous channel conditions [26].
VI-F3 Data Representation and Rate
Efficient data representation is a pivotal factor in achieving high data rates in deep learning AE-based communication systems. Traditional encoding schemes, such as one-hot vectors, exhibit inherently low spectral efficiency due to their sparse structure, which limits the amount of information that can be transmitted per symbol [100]. This inefficiency necessitates the exploration of alternative representation methods capable of enhancing throughput without incurring excessive computational costs. Furthermore, scaling these representations to accommodate large message dimensions becomes increasingly impractical, as it imposes significant demands on memory and processing resources [79]. Consequently, the limitations of current AE-based frameworks underscore the pressing need for innovative transmission schemes that can reconcile high data rates with practical implementation constraints [100].
To address these challenges, recent efforts have introduced new transmission paradigms specifically designed to increase data rates in neural communication systems, such as generalized data representation (GDR) [100]. GDR is effective in representing messages with large dimensions and has low complexity.
VI-F4 Data Scarcity
Data scarcity presents significant challenges in communication systems, particularly concerning training data for AE. Collecting extensive RF signal datasets is difficult, especially under diverse operational conditions and low SNR environments [102]. This difficulty leads to constraints for traditional signal classification methods due to limited training data [103]. Small sample sizes may lead to overfitting issues when training models [74]. This highlights a need for better handling of limited data. Challenges also exist in classifying wireless RF signals in complex environments with limited labeled training data, and progress on environment-independent AE deep learning models remains slow, which requires large labeled datasets for training. Achieving efficient communication with limited training data is also a discussed problem.
AE and their variants are explored as potential solutions. A VQ-VAE is proposed to augment training data for classifying wireless RF signals in complex environments with limited labeled training data [103]. Deep Denoising AE (DAEs) are highlighted as effective in scenarios with limited training data, which reportedly outperform existing solutions for channel modeling and precoding [138]. AEs are also demonstrated as feasible for joint communication and sensing of multiple targets, with a novel counting encoding method proposed that is particularly suitable for systems with a limited number of available snapshots [101]. A semi-supervised classification method combining SHAE and a semi-supervised linear discriminant is proposed to enhance generalization performance with limited labeled data [105]. A data-driven approach using AEs can maintain performance even when perfect channel information is unavailable [87].
Furthermore, the discussion around adapting to channel characteristics with measured data [63], optimizing training efficiency [19], and applying learning strategies without continuous on-device learning [110] implies the potential benefit of leveraging knowledge gained from one scenario (where data might be more abundant) to another (where data might be scarce). This aligns with the general concept of transfer learning, where a model trained on a related task or dataset is adapted for a new task with potentially less data. This particular aspect necessitates further deep investigation.
However, general challenges remain for AE, such as difficulties training with high code dimensions, which can lead to local minima, and the instability of some mutual information estimators with larger values, although the direct link between these general issues and limited data is not explicitly detailed [74]. The performance of AEs is also observed to be influenced by training procedures, architecture design, and hyper-parameter tuning, factors that can be complicated by limited data [24, 62]. The high complexity of end-to-end learning communication systems using AEs hinders further development, which requires significant computational resources, a challenge potentially exacerbated with complex limited datasets. The counting encoding method, used for limited snapshots, may lead to a lower detection rate compared to one-hot encoding. Performance may degrade significantly when domain regularities are not applicable, which could be an issue if limited datasets do not capture diverse real-world scenarios [16, 101]. This concludes the need for better handling of limited data by investigating the other existing methods (which could include some AE applications) as proposed in [139].
VI-F5 Emerging Hardware Technologies
AE-based approaches are increasingly aligning with next-generation communication hardware technologies, particularly neuromorphic computing and photonic neural networks. This convergence represents a paradigm shift towards energy-efficient, high-speed, and intelligent communication systems that can meet the demands of 6G and beyond. The neuromorphic joint source-channel coding (NeuroJSCC) framework represents a pioneering all-spike solution that combines neuromorphic sensing using event-driven cameras (DVS), spiking neural networks (SNNs) for encoding/decoding Impulse Radio (IR) for wireless transmission [140]. This system achieves remarkable energy efficiency with consumption proportional to spike activity, enabling picojoule-per-spike operation. However, End-to-end training of all components, including the SNNs and channel model, can be computationally demanding, where training separate models at different SNR levels is expensive. In addition, the complexity of the network structure might lead to over parametrization and overfitting, especially when increasing transmission rates does not necessarily lead to better performance.
A photonic encoder-decoder (PED) based on an all-optical variational autoencoder is a remarkable system using linear diffractive processors, photonic VAEs for high-throughput image transmission and WDM-based broadcast-and-weight architectures [141]. This is a clever implementation, since the system leverages optical signals to perform encoding, encryption, compression, and decoding without converting signals into electronic form and back (i.e., no analog-digital conversions). These frameworks demonstrate the potential for tera-operations per second (TOPS) throughput, far exceeding conventional digital signal processing capabilities. However, the experimental setup underlines potential challenges. The reliance on precise alignment of optical components in the adaptive training section suggests that even minor misalignments may affect performance. Although the adaptive tuning strategy is beneficial for handling variations in the system, the need for retraining and fine-tuning, particularly with a limited dataset, may limit rapid deployment and adaptability in dynamically changing transmission conditions.
The development of neuromorphic and photonic technologies is still in its early stages, with ongoing research needed to address material challenges and improve the maturity level of these technologies [142]. Despite these challenges, the potential benefits of integrating AE-based approaches with next-generation communication hardware are substantial, offering a transformative impact on the field of artificial intelligence and communication systems.
VII (Non-) and Differentiable Channel and Components
A differentiable channel refers to a channel model that can be represented mathematically and whose behavior can be differentiated with respect to its input. Specifically, it means that the channel output (received signal) can be expressed as a differentiable function of its input (transmitted signal). This property is crucial for training end-to-end neural network-based AE learning, where a single neural network represents the entire communication system (transmitter, channel, and receiver). The AE aims to learn an efficient representation of transmitted information, and having a differentiable channel allows gradients to flow through the entire system during training, as shown in Fig. 12.
VII-A Determining Differentiability
Whether a channel is differentiable depends on the specific characteristics of the channel model. a differentiable function (channel) typically has a tangent line at every point along its graph. If the tangent exists, the function is continuous, and its derivative exists, making it differentiable [143]. Here are some considerations:
Mathematical formulation
If the channel can be described by a mathematical function, such as linear channel, AWGN, or Rayleigh fading, it is likely to be differentiable. For example, the AWGN channel can be expressed as $y=x+n$, where $y$ is the received signal, $x$ is the transmitted signal, and $n$ is the noise. This channel is differentiable because its behavior can be expressed as a continuous function.
Discrete Effects
Non-differentiability often arises from discrete effects, such as quantization and modulation. If the channel introduces discrete operations, such as a threshold, it may not be differentiable. In such cases, alternative training methods such as reinforcement learning are needed.
Noise Models
Noise models matter. AWGN is differentiable due to its Gaussian noise, but other noise models, such as impulsive noise, may not be.
Physical limitations
Real-world channels have physical constraints, such as power limits or bandwidth limits. These constraints affect differentiability. For instance, clipping can make a channel non-differentiable.
Non-differentiableChannel SolutionsAlternatingtraining algorithmLearning stochasticchannel modelTwo-phase training strategyReinforcementlearning-based approachesGenetic algorithmsGradient-freetraining methodDifferentiableDSP algorithmsStochastic PerturbationTechniqueConditional GAN
Figure 13: Solutions for Non-differentiable Channel
VII-B Solution Approaches for Non-differentiable
Alternating training algorithm
The idea behind this is to optimize the transmitter and receiver independently. The iterative optimization process involves alternating between optimizing the receiver with a fixed transmitter and subsequently optimizing the transmitter with a fixed receiver. This approach aims to enhance the overall system performance through this iterative refinement. This can be achieved through an alternating optimization approach, where the receiver is trained using supervised learning while the transmitter is optimized using reinforcement learning [144, 145].
Differentiable DSP algorithms
An end-to-end differentiable AE channel require all components, including DSP algorithms, to be differentiable to enable gradient-based optimization. Traditional DSP algorithms often incorporate non-differentiable operations, hindering their direct integration into the differentiable channel. Therefore the authors of [114] proposed to modify blind phase search (BPS) algorithm and make it differentiable to include it in the end-to-end constellation shaping.
Learning stochastic channel model
This approach involves learning a surrogate channel model, either through supervised learning [146] or adversarial methods [147]. The key idea is to learn a differentiable generative model of the channel in the form of a GAN, which can then be used to train the AE, and to approximate end-to-end channel responses. The learned model is subsequently employed to train the transceiver. This enables direct optimization of the modulation and encoding components within an end-to-end AE model [148]. The accuracy of this learned model is crucial for the overall system performance. In [149], a channel-sensitive AE (CSAE) incorporates a pre-trained conditional GAN (CGAN) to model optical fiber communication systems with varying impairments.
Two-phase training strategy [26]
With this approach, first train the AE using a channel model designed closely to the real channel such as a stochastic channel model. Upon deploying the trained model for real-wold transmission, performance is contingent on the model’s accuracy. To mitigate this, based on measurement data, a find-tuning phase focusing exclusively on the receiver part of the AE, which leads to a sub-optimal training.
Reinforcement learning-based approaches
The idea is to train an neural network (NN)-based transmitter using policy gradients while employing a non-differentiable receiver that treats detection as a clustering problem.
Stochastic perturbation techniques
A gradient free optimization techniques called simultaneous perturbation stochastic optimization has been proposed in [150]. It is a novel, single-stage approach to train deep learning-based communication systems directly in real-world channel environments. By circumventing the need for explicit channel modeling, this way leverages stochastic approximation to estimate gradients, enabling end-to-end optimization.
Gradient-free training method
It is a derivative-free optimization method to eliminate the need for a differentiable channel model in AE training. This enables simultaneous optimization of the encoder and decoder for arbitrary, even non-numerical, channel environments. The authors of [151] proposed the use of the cubature kalman filter (CKF) which offers gradient-free training while maintaining accuracy, thereby facilitating online adaptation.
Genetic algorithms (GAs)
While traditional end-to-end learning relies heavily on backpropagation, which requires the system to be differentiable, genetic algorithms offer an alternative optimization method that does not rely on gradient information [152, 153, 154, 155]. The advantage of using GAs are as (1) optimization without gradients: GAs optimize a population of solutions through evolutionary processes such as selection, crossover, and mutation. This optimization process does not require the computation of gradients and can be applied to non-differentiable functions. Thus, if the channel in the communication system is non-differentiable, due to involving quantization, discrete noise, or other non-continuous processes, traditional gradient-based methods fail to optimize the AE effectively. GAs, can still optimize the system by treating it as a black-box and relying on fitness evaluations rather than gradient information, (2) Fitness evaluation: In the context of an AE, GAs can evaluate the fitness of each individual in the population based on how well the AE reconstructs the input data after it has passed through the channel. The fitness function can be designed to measure the reconstruction error or any other relevant performance metric, and (3) Population diversity: GAs maintain a population of potential solutions, which can help to explore the solution space more thoroughly compared to gradient descent methods that could get stuck in local minima. This can be particularly beneficial in complex, non-differentiable landscapes.
Conditional GAN
The conceptual structure involves the use of a conditional GAN to represent the conditional distribution of a channel [156]. This allows channel effects to be learned directly from the data rather than relying on expert knowledge or assumed mathematical models. The learned conditional GAN acts as a surrogate for the real channel, which allows gradients to pass through to the transmitter, thus enabling the back-propagation of gradients to the transmitter even when the real channel is unknown, which is facilitating end-to-end learning. For time-varying channels, received pilot data is also included as part of the conditioning information. This enables the conditional GAN to generate data corresponding to specific instantaneous channel conditions.
VIII Lessons Learned
Based on the exhaustive review described above, a compilation of recommended methodologies has been identified to ensure robust design of AE and reliable deployment in practical settings.
Regularization Strategies
The physical constraints such as power limits and computational efficiency such as sparsity must be balanced. In wireless communication, it shall use constraint in the encoder to enforce physical limits such as power normalization. Dropout is less common; instead, GAN-based adversarial training helps model channel variations. In addition, batch normalization is widely utilized instead of dropout. In optical communication, it has been seen that sparse AEs can reduce complexity for long-haul dispersion compensation, and VAE-based regularization (KL divergence) improves robustness to phase noise.
Loss Function Choices
for Regression Tasks (symbol recovery), MSE dominates but fails at low SNR; Huber loss is more robust. For Classification Tasks (modulation detection), MI-based losses excel in capacity-approaching codes, while CCE is preferred for modulation tasks. Hybrid losses have also been incorporated such as combine MSE + PAPR penalty for OFDM waveform design.
Hyperparameter Tuning Techniques
Like activation functions (ELU/Tanh) and learning rates (cyclic) are critical for convergence in noisy environments. LeakyReLU/ELU is preferred over ReLU for gradient flow in fading channels, while Tanh can mitigate exploding gradients in optical systems. Furthermore, cyclic LR can avoid vanishing gradients. In terms of learning rate, cyclic learning rates help escape local minima in end-to-end training.
Deployment Reliability
In the context of hardware constraint deployment, quantization-aware training such as 1-bit AE can reduce the FPGA resource. In addition, hybrid model-based AE designs facilitate FPGA implementation. Transfer learning can fine-tune AEs on real-world data and helps in channel mismatch mitigation.
IX Future Research Areas
Future research can further advance the application of AEs in communication systems. Therefore, there are numerous unresolved challenges and inherent limitations when it comes to using AEs in communication systems. It is imperative that these issues undergo comprehensive investigation to uncover novel solutions or refine existing methodologies that address these challenges.
IX-1 Advanced AE architectures